# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we are first able to partition transitive triangles into interior neighbors, exterior neighbors, and back arcs. We also partition the nodes of a graph based on distance from the minimum out-degree node. These operations establish the Graph Level Order (GLOVER) data structure. This establishes a well-ordering of neighborhoods, and allows for the construction of a decreasing sequence of subsets of nodes. This sequence shows the non-existence of counterexamples to the SSNC and precisely identifies the required node. Further, our approach finds the occurrence of dense graphs inside neighborhoods. This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property. Beyond theoretical implications, the algorithm and data structure have practical applications in data science, network optimization and algorithm design.
Keywords: data structure, proof by contradiction, rooted neighborhood, exterior neighbor, minimum degree node, transitive triangle, load balancing
@@ 1 Introduction
Unlike previous approaches, this framework allows us to locate vertices that satisfy the SSNC explicitly by way of a path to those vertices. The development of this data structure has been an interdisciplinary challenge, which drew insights from combinatorics, algorithm design, and network theory. With a background in computer programming and data science, though, our search was for not only provable theorems, but usable and verifiable algorithms. Even the abstract concepts that we produced via group theory, we have tried to maintain algorithmic accessibility throughout.
What follows is a journey from intuition to implementation. We develop and formalize a conjecture, support it with rigorous mathematical claims, and trace computational realizations. This paper aims to offer both clarity and depth while grounding abstract theory in practical, verifiable methods. Developing the GLOVER data structure presented a significant challenge, requiring insights from various areas of mathematics and computer science. Along the way, we will make rigorous mathematical claims, justify them thoroughly, and explain the motivations behind key decisions. This paper aims to provide both clarity and depth in tackling one of graph theory’s most enduring puzzles.
@@
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second out-neighborhoods of $v$, respectively.
This conjecture was first published by Dean and Latka in 1995 [7], where they proposed a related version specific to tournaments, which are complete oriented graphs. The tournament case was proved by Fisher [8] in 1996, with an alternative proof later provided by Havet and Thomassé [9] in 2000.
The work on the solutions to the Dean Tournament Conjecture is very important in the study of second out-neighborhood conjecture. Fisher’s approach [8] did not yield a method for locating a specific vertex, but it did prove that such a vertex must exist. The method did so with probability, by showing that there exists at least one node whose second out-neighborhood must be at least as large as its first. Later, Havet and Thomassé [9] advanced the problem further by using median orders to provide a more constructive approach. This technique identifies a candidate vertex and a path to that vertex. The method relies on concepts related to median orders, which is an NP-hard problem and introduces computational complexities to their approach. Still, this was a significant breakthrough in combinatorics, particularly for tournament theory, as median-orders were now available for usage.
Kaneko and Locke [10] showed that the SSNC holds for graphs with a minimum degree of at most six. This work provided a ground level for a lot of the research on the SSNC problem. It also brought in another NP-Hard problem, dominating sets, into the realm of the SSNC. Unfortunately, for cases greater than six, the complexity was too great to solve efficiently.
@@
These results originate from Constraint Satisfaction Programming (CSP). This is a method that tries to frame the SSNC as a series of logical constraints to be satisfied and optimized by way of SAT solvers. This direction has led to increasingly tighter bounds, but it operates within the realm of NP-hard problems, which makes both scalability and generalization a challenge. Also, resolving SSNC through CSP offers little insight into the graphs themselves, which is an issue we aim to resolve with our constructive algorithmic approach.
@@
In recent work, Diaz et al. [6] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [3], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC. The study by Diaz et al. examines randomly generated graphs, specifically using binomial random graphs with random orientations. It demonstrates that in these constructions there are typically Seymour vertices. These findings are valuable for understanding the average-case behavior of oriented graphs, and this is a major step towards resolving the conjecture. However, identifying Seymour vertices in randomly generated cases fails to identify properties of graphs that either guarantee or preclude the conjecture’s truth. There are many fundamental set-theoretic and graph-theoretic questions that remain open. These are questions that concern the connectivity of graphs, degree distribution, and neighborhood growth.
@@
Nevertheless, the Diaz et al. paper does make meaningful progress by considering the contrapositive of the Seymour conjecture. In particular, Propositions 4 and 5 in the paper explore the implications of the conjecture being false. They show that in such a case, there would be infinitely many strongly connected oriented graphs without Seymour vertices, even if we had a bounded minimum out-degree. This line of reasoning draws resemblance to the Decreasing Neighborhood Sequence Property (DNSP) introduced in this paper. Though not mentioned in their work and having different implications, the counterexample modeling echoes the motivations that led to the development of DNSP here.
@@
Building upon this understanding of local structures, our work examines a particularly insightful contribution that comes from Brantner et al. [2], who examined the role of specific subgraph structures in relation to the Seymour Second Neighborhood Conjecture. Their work introduced and formalized two key motifs: transitive triangles and Seymour diamonds in the context of SSNC. A transitive triangle consists of three nodes and three directed edges, forming a configuration where two edges share a common head—an arrangement that creates a kind of shortcut in the graph’s reachability. A Seymour diamond involves four directed edges connecting four nodes, with two edges originating from a common source and converging again at a shared target. Brantner et al. showed that if a directed graph is free of transitive triangles, then it necessarily contains a Seymour vertex—a node whose second out-neighborhood is at least as large as its first. This result not only builds intuition about the conjecture but also suggests that the presence of certain local configurations might be obstructive to the existence of a Seymour vertex. Our work examines how larger, layered neighborhood interactions—particularly in the presence of overlapping or nested substructures—can be systematically analyzed through a new data structure.
The concept of $m$-free graphs is particularly relevant to this paper[1]. A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$. In his paper, Daamouch works with the concept of induced subgraphs and explores their role in the degree and properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
@@
Variations of the question have asked whether the conjecture holds for graphs lacking certain substructures. For example, Fidler and Yuster [19] proved the conjecture true for tournaments when multiple edges are removed from the same node and orientations of complete graphs missing a star. Later [17] proved the conjecture true for oriented graphs missing $C_{4},\overline{C_{4}},S_{3}$ chair or co-chair. Following this, Ghazal proved [20] the conjecture true for tournaments missing a star. Daamouch et al. [18] proved the conjecture true for tournaments missing two stars or disjoint paths .
@@
Also of note is the 2021 poster presentation by Illia Nalyvaiko [21], which—despite its informal venue—offered another contrapositive-style approach to explore the SSNC along lines somewhat similar to the Decreasing Neighborhood Sequence Property (DNSP). In his poster, he introduced a penalty function to evaluate the likelihood of a vertex of a counterexample. While this is a different intuition than DNSP, both approaches represent a shift from verifying the conjecture to questioning the possibility of the bounds of the conjecture itself. Given the historical significance of this problem, Nalyvaiko’s contrapositive-style approach, while informal, offered a novel perspective that resonated with the motivations behind our Decreasing Neighborhood Sequence Property.
@@
Rather than simply asking whether a graph satisfies the conjecture by possessing a Seymour vertex, we turn our attention to the nodes that fail to be Seymour vertices. We ask this question not explicitly for a particular node, but for every node in the graph. Doing it this way enables us to engage in a search, because we are looking for conditions that need to be satisfied for every node in the graph. As such, this proof for a degree doubling node is like building a house of cards. There will be a lot of things that need to be true for every node to fail to be a degree doubling node. As we continue to try to build this house of cards, what we will see is that it will lead to one of two things. Either it eventually collapses, leading to a contradiction and producing that degree doubling node. The other possibility is that we finish constructing the house, and every node fails to be a degree doubling node. This would prove the conjecture false.
@@
Beyond its theoretical appeal, the SSNC has practical applications in domains like social networks, epidemiology, algorithm design, and the network A/B testing. For example, in disease control, identify individuals whose influence spans two levels of contact target interventions more effectively and mitigate outbreaks. In network-based A/B testing, it is crucial to have disjoint treatment and control groups to avoid interference. That is a fundamental feature of the GLOVER data structure. Furthermore, within these communities, the SSNC helps us identify highly influential individuals, allowing for more strategic experimental design and more accurate evaluation of interventions.
@@
Unlike traditional theoretic graph theory proof approaches, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.

### v12

@@ Abstract
The Seymour Second Neighborhood Conjecture (SSNC) claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we establish the Graph Level Order (GLOVER) data structure, which orders the nodes by shortest path from a minimum out-degree node and establishes a well-ordering of rooted neighborhoods. This data structure allows for the construction of decreasing sequences of subsets of nodes and allows us to partition transitive triangles into distinct sets. The decreasing sequence of nodes shows the non-existence of counterexamples to the SSNC and precisely identifies a path to the required node. Further, our algorithmic approach finds the occurrence of dense graphs inside the rooted neighborhoods. Beyond theoretical implications, the algorithm and data structure have practical applications in data science, network optimization and algorithm design.
Keywords: data structure, decreasing neighborhood sequence property, rooted neighborhood, exterior neighbor, minimum degree node, transitive triangle, load balancing
@@ 1 Introduction
Unlike previous approaches, this framework allows us to locate vertices that satisfy the SSNC explicitly by way of a path to those vertices. The development of this data structure has been an interdisciplinary challenge, which drew insights from combinatorics, algorithm design, and network theory. With a background in computer programming and data science, our search was for not only provable theorems, but usable and verifiable algorithms. Even with the abstract concepts we produced via group theory, we have tried to maintain algorithmic accessibility throughout.
@@
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G$, there exists a vertex $v\in G$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second out-neighborhoods of $v$, respectively.
This conjecture was first published by Nathaniel Dean and Brenda Latka in 1995 [8], where they proposed a related version specific to tournaments, which are complete oriented graphs. The tournament case was proved by Fisher [11] in 1996, with an alternative proof later provided by Havet and Thomassé [18] in 2000.
The work on the solutions to the Dean Tournament Conjecture is very important in the study of the second out-neighborhood conjecture. Fisher’s approach [11] did not yield a method for locating a specific vertex, but it did prove that such a vertex must exist. The method did so with probability, by showing that there exists at least one node whose second out-neighborhood must be at least as large as its first. Later, Havet and Thomassé [18] advanced the problem further by using median orders to provide a more constructive approach. This technique identifies a candidate vertex and a path to that vertex. Median orders is an NP-hard problem and introduces computational complexities to their approach. Still, this was a significant breakthrough in combinatorics, particularly for tournament theory, as median-orders were now available for usage.
Kaneko and Locke [20] showed that the SSNC holds for graphs with a minimum degree of at most six. This work provided a ground level for a lot of the research on the SSNC problem. However, it also brought another NP-Hard problem, dominating sets, into the realm of the SSNC. Unfortunately, for cases greater than six, the complexity was too great to solve efficiently.
@@
These results originate from Constraint Satisfaction Programming (CSP). This is a method that tries to frame the SSNC as a series of logical constraints to be satisfied and optimized by way of SAT solvers. This direction has led to increasingly tighter bounds, but it again operates within the realm of the NP-hard problem CSP. This makes both scalability and generalization a challenge. Also, resolving SSNC through CSP offers little insight into the graphs themselves, which is an issue we aim to resolve with our constructive algorithmic approach.
@@
In recent work, Diaz et al. [9] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [1], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC. The study by Diaz et al. examines randomly generated graphs, specifically using binomial random graphs with random orientations. It demonstrates that in these constructions there are typically Seymour vertices. These findings are valuable for understanding the average-case behavior of oriented graphs, and this is a major step towards resolving the conjecture. However, identifying Seymour vertices in randomly generated cases fails to identify properties of graphs that either guarantee or prevent the conjecture’s truth. There are many fundamental set-theoretic and graph-theoretic questions that remain open. These are questions that concern the connectivity of graphs, degree distribution, and neighborhood growth.
@@
Nevertheless, the Diaz et al. [9] paper does make meaningful progress by considering the contrapositive of the Seymour conjecture. In particular, Propositions 4 and 5 in the paper explore the implications of the conjecture being false. They show that in such a case, there would be infinitely many strongly connected oriented graphs without Seymour vertices, even if we had a bounded minimum out-degree. This line of reasoning draws resemblance to the Decreasing Neighborhood Sequence Property (DNSP) introduced in this paper. Though not mentioned in their work and having different implications, the counterexample modeling echoes the motivations that led to the development of DNSP here.
@@
Building upon this understanding of local structures, our work examines a particularly insightful contribution that comes from Brantner et al. [2], who examined the role of specific subgraph structures in relation to the SSNC. Their work introduced and formalized two key patterns: transitive triangles and Seymour diamonds in the context of SSNC. A transitive triangle consists of three nodes and three directed edges, forming a configuration where two edges share a common head—an arrangement that creates a shortcut in the graph’s reachability. A Seymour diamond involves four directed edges connecting four nodes, with two edges originating from a common source and converging again at a shared target. Brantner et al. showed that if a directed graph is free of transitive triangles, then it necessarily contains a Seymour vertex—a node whose second out-neighborhood is at least as large as its first. This result not only builds intuition about the conjecture but also suggests that the presence of certain local configurations might be obstructive to the existence of a Seymour vertex. Our work examines how larger, layered neighborhood interactions—particularly in the presence of overlapping or nested transitive triangle substructures—can be systematically analyzed through a new data structure.
The concepts of $m$-free, $k$-transitive, and $k$-anti transitive graphs are also relevant to this paper [6]. A directed graph $D$ is called $m$-free if it contains no directed cycles of length at most $m$. Daamouch also considered $k$-transitive graphs, which are graphs that had these transitive shortcuts for $k\leq 6$, and $k$-anti transitive graphs, for $k\leq 4$. These are graphs without these shortcuts. Daamouch related these extensions of transitivity and anti-transitivity to graphs necessarily having a Seymour vertex. Later, [16] Hassan et al. extended these results to 6-anti transitive graphs.
Transitive triangles are fundamental in the structure of real-world networks, such as those underlying social media and friend recommendation systems. The SSNC is not just a mathematical problem, we are attempting to solve a real world problem. Therefore, any algorithm or framework attempting to resolve the SSNC must be capable of operating meaningfully on graphs that contain them.
The Graph Level Order data structure provides exactly this capability. By organizing nodes based on their distance from a minimum out-degree vertex, it reveals that transitive triangles are not uniform obstacles. Instead, they can be partitioned into distinct cases based on the relative distance of their vertices. This insight is not simply novel, it is necessary. Without formalizing these rooted neighborhoods, critical properties such as decreasing neighborhood size and the influence of back arcs would remain unprovable. The Graph Level Order enables the distinction and handling of cases that were previously inseparable, and is thus essential to the approach developed in this paper.
@@
Variations of the question have asked whether the conjecture holds for graphs lacking certain substructures. For example, Fidler and Yuster [12] proved the conjecture true for tournaments when multiple edges are removed from the same node and orientations of complete graphs missing a star. Later Mniny and Ghazal [21] proved the conjecture true for oriented graphs missing $C_{4},\overline{C_{4}},S_{3}$ chair or co-chair. Following this, Ghazal proved [14] the conjecture true for tournaments missing a star. Daamouch et al. [7] proved the conjecture true for tournaments missing two stars or disjoint paths.
@@
Also of note is the 2021 poster presentation by Illia Nalyvaiko [22], which offered another contra-positive-style approach to explore the SSNC somewhat similar to the Decreasing Neighborhood Sequence Property (DNSP). In his poster, he introduced a penalty function to evaluate the likelihood of a vertex of a counterexample. While this is a different intuition than DNSP, both approaches represent a shift from verifying the conjecture to questioning the possibility of the bounds of the conjecture itself. Given the historical significance of this problem, Nalyvaiko’s contrapositive-style approach offered a novel perspective that resonated with the motivations behind our Decreasing Neighborhood Sequence Property.
@@
Rather than simply asking whether a graph satisfies the conjecture by possessing a Seymour vertex, we turn our attention to the nodes that fail to be Seymour vertices. We ask this question not explicitly for a particular node, but for every node in the graph. This enables us to engage in a search, because we are looking for conditions that need to be satisfied for every node in the graph. As such, this proof for a degree doubling node is like building a house of cards. There will be a lot of information that needs to be true for every node to fail to be a degree doubling node. As we continue to try to build this house of cards, what we will see is that it will lead to one of two things: either the house will eventually collapse, leading to a contradiction and producing that degree doubling node, or we finish constructing the house, and every node fails to be a degree doubling node. This later statement would prove the conjecture false.
@@
Beyond its theoretical appeal, the SSNC has practical applications in domains like social networks, epidemiology, algorithm design, and network A/B testing. For example, in disease control, we need to identify individuals whose influence spans two levels of contact. We could target interventions more effectively and mitigate outbreaks. In network-based A/B testing, it is crucial to have disjoint treatment and control groups to avoid interference. That is a fundamental feature of the Graph Level Order data structure. Furthermore, within these communities, the algorithm for SSNC helps us identify highly influential individuals, allowing for more strategic experimental design and more accurate evaluation of interventions.
@@
Unlike the purely existential proof approaches we have considered, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
This paper claims a complete proof of the SSNC using a proof by contradiction-based argument, fundamentally supported by a novel and essential data structure: the Graph Level Order. While the idea of partitioning nodes into neighborhoods is a common technique. Our central findings become evident precisely where BFS ends. We formalize the intricate relationships between neighborhoods, within themselves, to each other, and with the root node. This establishes a powerful framework for understanding the SSNC. The Graph Level Order is not merely a traversal algorithm. It systematically defines the necessary implications for these neighborhoods to uphold the conjecture. For instance, it allows us to decompose transitive triangles, a long-standing barrier to progress on the conjecture, into six distinct, manageable cases. Our approach is constructive and grounded in standard algorithmic tools like partition, cycles, divide and conquer, and traversal and offers linear-time complexity. Beyond merely solving the conjecture, it provides a straightforward blueprint for identifying the node that satisfies it. All key lemmas, theorems and corollaries are all rigorously proven and all terminology strictly adheres to standard graph-theoretic conventions.